OpenStudy (anonymous):

a copy machine makes 207 copies in 5 minutes and 45 seconds, how many copies does it make per minute?

jimthompson5910 (jim_thompson5910):

5 minutes and 45 seconds = 5.75 minutes

jimthompson5910 (jim_thompson5910):

(207 copies)/(5.75 minutes) = (207/5.75) copies per minute (207 copies)/(5.75 minutes) = 36 copies per minute